# Efficient inference of quantum system parameters by Approximate Bayesian Computation

### v2

@@ V.2 Two-level atom
V.2.1 Summary statistic
@@
Figure 7: ABC performance for the linear optomechanical system using the total time to emit $N=50$ photons as the summary statistic, for two different values of detuning: $\Delta=-1\kappa$ (a-c) and $\Delta=-7\kappa$ (d-f). In each plot the ABC-based posterior distributions (different colours for different thresholds $\epsilon$) are presented against the true posterior (black). As the latter is obtained by performing brute-force numerics, this precludes from presenting the performance ABC averaged over different photon-click trajectories. That is why, the insets present (in contrast to Fig. 6) the fidelity (46) averaged over 10 implementations of ABC for each threshold, while in each of the six cases a different photon-click trajectory $D_{t}$ is considered. We find good agreement with the true posterior in all cases, demonstrating the validity of ABC. Moreover, the performance is satisfactory independently of detuning value—being in line with our understanding of photon-click statistics and the coefficient of variation $\gamma$ presented in Fig. 5(b).
V.2.2 Results
@@
In the insets of Fig. 6 we calculate the fidelity (46) between the posteriors generated through ABC and the respective true ones. We average over 20 different trajectories (i.e. true posteriors) and 10 unique implementations of ABC for each trajectory, to show the average performance of ABC. The number of samples required to saturate the performance is now larger though, so successful implementation requires longer computation time.
Figure 8: ABC performance for the non-linear optomechanical system using the total time to emit $N=80$ photons as the summary statistic, for two different values of detuning: $\Delta=\Delta_{n=1}\approx-2.8\kappa$ (a-c) and $\Delta=\Delta_{n=2}\approx-5.6\kappa$ (d-f). In each subplot (a-f) a different photon-click trajectory is considered and, as before, ABC-based posterior distributions (different colours for different thresholds $\epsilon$) are presented against the true posterior (black). In (a)-(c) we see a reasonable performance of ABC, though clearly less accurate than in the case of linear system. In (d)-(f), however, the algorithm fails to reproduce the posterior in all the cases, resulting in low fidelity (averaged over 10 ABC implementations) within each inset. As expected from Fig. 5(c), the total time $t_{N}$ can now no longer be considered as a sufficient summary statistic, especially around $\Delta=\Delta_{n=2}$.
@@ V.3 Linear optomechanical system
V.3.1 Summary statistics
In Sec. IV.3.2, it was shown that the intensity of emitted light, i.e. the exponent of the exponential waiting-time distribution, should be almost completely characteristic of the emission profile for a linear optomechanical system. Hence, the total time for a fixed number of photon emissions, $t_{N}$ in Eq. (42), should again be sufficient as a summary statistic. However, as the simulation of dynamics is more computationally exhaustive now, we restrict ourselves to 50 photons, i.e. $S(D_{t})=t_{N=50}(D_{t})$. This is (at least approximately) sufficient due to the simplicity of the exponential average waiting-time distribution—assured by the (almost) Poissonian emission profile.
V.3.2 Results
@@
We again quantify the performance of the algorithm through the fidelity (46). However, unlike the two-level atom, obtaining true posteriors requires numerical solutions to the dynamics, which are extremely time consuming. We instead show the results of three indicative trajectories for $\Delta=-1\kappa$ in Fig. 7(a)-(c) and for $\Delta=-7\kappa$ in Fig. 7(d)-(f), while the insets now show just the average fidelity over 10 implementations of the ABC algorithm—being no longer averaged over different photon-click trajectories. Nevertheless, we see reasonable agreement for all the trajectories considered, suggesting the performance is relatively consistent.
In particular, we also see consistent performance for the different values of detuning. This is in agreement with expectations from the coefficient of variation $\gamma$ presented in the inset of Fig. 5(b), which is mostly invariant to changes of the detuning. Moreover, as the statistics form an exponential distribution, the emitted light is almost coherent with no inter-photon correlations, $\forall\tau:\;g^{(2)}(\tau)\approx 1$ in Fig. 4, meaning that the total time for a fixed number of emissions (42) is highly informative. We therefore see an improved performance compared to the two-level atom.
@@ V.4 Non-linear optomechanical system
Finally, we consider the non-linear optomechanical system. As already emphasised, the dynamics is now complex, resulting in non-trivial photon statistics. Nevertheless, we first proceed with the same summary statistics approach to the ABC algorithm as in the previous two cases, in order to explicitly demonstrate that, as expected, it will not be satisfactory.
V.4.1 Summary statistics
@@
V.4.2 Results
We show the results compared to true posteriors obtained by numerical means in Fig. 8, as for the linear optomechanical system. For Fig. 8(a)-(c) we set $\Delta=\Delta_{n=1}$, while for Fig. 8(d)-(f) we set $\Delta=\Delta_{n=2}$. As expected, in the $n=1$ regime, we see a good performance of ABC, capable of reproducing the posterior to a high degree of accuracy. For the $n=2$ regime, however, the algorithm does not reproduce the posterior, being capable only to eliminate the “inferable” values of the detuning around $\Delta_{n=1}$ from the distribution.
@@ V.4.2 Results
Figure 9: Histogram of waiting times $\tau$ in-between subsequent photon-clicks for detuning $\Delta_{n=1}$ in (a) and $\Delta_{n=2}$ in (b). In each case, the data represents average occupancies for 2000 trajectories, each of $N=80$ photo-clicks. By focussing on the short-time regime in the above ($\tau\leq 8\kappa^{-1}$), we see a clear difference between the distributions: (a) – an exponential decreasing modestly due to uncorrelated single-photon ($n=1$) emissions being dominant; versus (b) – a distribution dropping steeply at low $\tau$ due to double-photon ($n=2$) emissions being favoured. The error bars mark the standard deviation of the trajectory-samples used.
@@
Figure 10: ABC performance for the non-linear optomechanical system employing the distanced-based full data approach. In particular, the acceptance criterion of the ABC utilises the L2-norm (47) between the histograms representing the waiting-time distribution in the range of short-time intervals, with each trajectory consisting of $N=80$ photon-clicks. As in Fig. 8, two different values of detuning are considered: $\Delta=\Delta_{n=1}\approx-2.8\kappa$ (a-c) and $\Delta=\Delta_{n=2}\approx-5.6\kappa$ (d-f), with each subplot (a-f) being computed for a different photon-click trajectory. Importantly, in all the cases the ABC algorithms is now capable of reproducing the true posterior (black), while a less stringent $\epsilon$-threshold (different colours) must now be chosen due to much higher numbers of samples, $\nu$, being required for the method to converge.
V.5 Non-linear optomechanical system with the distance-based full data approach to ABC
In order to adapt ABC to the case of a non-linear optomechanical system, we could try to extend the number of summary statistics used within the algorithm, so that extra ones are included that, e.g., distinguish distinct types of inter-photon correlations. Identifying these is a hard task, especially for a complex system with dynamics being far from stationary over a significant amount of time. Instead, it is known to often be easier and more effective for the ABC algorithm to simply use the full (detection) data and compare it against the samples stored in the library by resorting to some statistical distance [63].
For the non-linear optomechanical system, we ignore non-stationarity of the detection data and consider it to be accurately represented for a given trajectory by the histogram of waiting times in between photon-click events. In Fig. 9, we present the so-obtained waiting-time distributions built upon considering 2000 simulated trajectories, each one containing $N=80$ photon-clicks, for the two detuning values of interest: $\Delta_{n=1}$ and $\Delta_{n=2}$.
We see that there is a very different distribution for these two values of detuning, which should be distinguishable by considering some statistical distance—here, we consider the L2-norm between the histograms, i.e.:
$\delta(\zeta(D_{t}),\zeta(D^{\prime}_{t}))=$ $\sqrt{\sum\limits_{i}\left(x_{i}-x^{\prime}_{i}\right)^{2}}\,,$ (47)
where $\zeta(D_{t})$ is the time-binned histogram of the waiting times of the data, and the $x_{i}$’s are the respective sizes of histogram entries. Within the ABC algorithm, we now modify the acceptance rule in Eq. (7), so that it uses the distance between the histograms $\zeta$, rather than between their particular summary statistics $S$. With such a modification, we compute the ABC-based posterior distributions, which we present in Fig. 10.
@@ V.5 Non-linear optomechanical system with the distance-based full data approach to ABC
Crucially, the ABC algorithm is now capable to reproduce the true posterior in both $n=1$ and $n=2$ regimes. For the $n=1$ regime, we see in Fig. 10(a-c) similar levels of performance of the ABC as in Fig. 8(a-c). However, for $n=2$ we observe a significant improvement in Fig. 10(a-f). As we tighten the $\epsilon$-thresholds in Fig. 10, we observe that none of the samples is sometimes accepted—indicating that these would require a much higher number of samples than $\nu=4\times 10^{5}$ here used. Nevertheless, for looser thresholds we manage to obtain good results even with the number of samples at hand.
@@ VI Conclusions
In this paper, we have identified the computation of the likelihood function as a fundamental bottleneck in quantum inference problems that resort to Bayesian analysis and aim at reconstructing the posterior distribution. This is a consequence of a dramatic growth of the underlying Hilbert space with the system size in the quantum setting. We argue, however, that such a chokepoint can be, in principle, circumvented by resorting to approximate sampling-based methods that require only the measurement data to be simulated—in contrast to actually constructing the likelihood of observing a particular data. We show that the former can still be performed efficiently in the quantum setting by resorting to quantum Monte-Carlo techniques [50, 51], also in case of dissipative systems with some degrees of freedom being inaccessible. Upon building a sufficiently large library of measurement data, we demonstrate that likelihood-free methods—in particular, the Approximate Bayesian Computation (ABC) algorithm [58, 59]—can then be used to reconstruct posterior distributions in quantum inference tasks.
@@
In order to show the applicability of our methods, we consider examples of a two-level atom and an optomechanical system, both being driven by classical light and continuously probed by photodetection. The former allows us to study the trade-off between the accuracy and efficiency of the ABC algorithm, thanks to the inference problem being actually solvable analytically. In stark contrast, in the latter case the sampling-based method is often essential to obtain results within a reasonable computation time. Yet, the understanding of the quantum statistics exhibited by the emitted photons [73] (bunching/antibunching) is then crucial to choose appropriate statistical tests employed within the ABC algorithm that allow to distinguish well between different datasets.
@@
In each case we demonstrate explicitly how the ABC algorithm should be tailored, in particular, we identify the summary statistic of the measured data that is sufficient to characterise the underlying process. Moreover, we study how to choose the thresholds for the acceptance criteria used in rejection sampling, which, if chosen too tight, may require impractically large numbers of samples. For the non-linear optomechanical system we resort to the full data distance-based approach [63], which leads to reasonable results despite minimal calibration. However, let us note that if one (unlikely) requires to reconstruct the posterior in the full parameter range, then ABC can be used to infer an “intermediate” posterior that is already narrowed down to relevant parameter ranges based on the data. As a result, it can be used afterwards as a prior by more computationally demanding methods, which can then be already effective thanks to the restricted parameter range.
@@
As our formalism can in principle be applied straightforwardly to other types of open quantum systems, we believe that our results pave the way for using likelihood-free inference methods across different experimental platforms. A natural next step is to combine these in the quantum setting with other statistical inference techniques such as the Markov chain Monte Carlo methods and the Metropolis-Hastings algorithm, as well as particle filters [54]. These are widely used to avoid the need of considering the full parameter space that, in contrast to our work, may pose a serious challenge when inferring multiple (multidimensional) parameters from the measurement data. We leave this as an interesting line of research to be pursued in the future.

### v3

@@ V.2 Two-level atom
V.2.1 Summary statistic: total time
@@
Figure 7: ABC performance for the linear optomechanical system with total time to emit $N=50$ photons used as the summary statistic, for two different values of detuning: $\Delta=-1\kappa$ (a-c) and $\Delta=-7\kappa$ (d-f). In each plot the ABC-based posterior distributions (different colours for different thresholds $\epsilon$) are presented against the true posterior (black). As the latter is obtained by performing brute-force numerics, this precludes from presenting the performance ABC averaged over different photon-click trajectories. That is why, the insets present (in contrast to Fig. 6) the fidelity (46), $F$, for each of the six different photon-click trajectories $D_{t}$ considered (still averaged over 10 implementations of ABC for each threshold). We find good agreement with the true posterior in all cases, demonstrating the validity of ABC. Moreover, the performance is satisfactory independently of detuning value—being in line with our understanding of photon-click statistics: the total time and its signal-to-noise ratio (SNR) presented in Fig. 5(b).
Results
@@
In the insets of Fig. 6 we show the average fidelity (46) between the posteriors generated through ABC and the respective true ones. We average over 20 different trajectories (i.e. true posteriors) and 10 unique implementations of ABC for each trajectory, to show the average performance of ABC. We see clearly that the number of samples $\nu$ required to saturate the performance of ABC is greater for $\Delta=7$ than $\Delta=1$. What we also see, within the range of $\nu$ considered, is that the performance is saturated for all but the tightest $\epsilon$-threshold value. However, for $\Delta=7$ there is hardly any improvement in the maximal average fidelity attained as $\epsilon$ is reduced. This is a consequence of trajectories possessing a much longer total time—recall Fig. 5(a) where $\langle t_{N}\rangle\approx 500(10000)$ for $\Delta=1(7)$. Hence, as it is the difference between total times that is used for acceptance, for thresholds to have the same significance these should be rescaled by a factor of $\approx 20$. However, we must avoid any adjustments of the ABC algorithm depending on the $\Delta$-value, as these would require the estimated parameter to be known in advance. Hence, we are forced to select a value of $\epsilon$ that to the best extent is tight enough within the whole parameter range $\Delta\in[0,10]$ considered—for example, from Fig. 6 we may infer $\epsilon=20\kappa^{-1}$ as an ideal candidate for all-round performance (given $\nu=20\times 10^{4}$ samples).
Lastly, let us comment on the bias exhibited by the posteriors in Fig. 6, which naively seem to be shifted to the left(right) for $\Delta=1(7)$. This, however, is a standard artefact of Bayesian inference and inspecting performance based only on a single trajectory. Moreover, as our summary statistic based on the total time is not exactly sufficient, it may also introduce some systematic errors. Nevertheless, when multiple trajectories are analysed, e.g. to compute the average fidelities in Fig. 6, we observe that the bias is washed out on average.
Figure 8: ABC performance for the non-linear optomechanical system with total time to emit $N=80$ photons used as the summary statistic, for two different values of detuning: $\Delta=\Delta_{n=1}\approx-2.8\kappa$ (a-c) and $\Delta=\Delta_{n=2}\approx-5.6\kappa$ (d-f). In each subplot (a-f) a different photon-click trajectory is considered and, as before, ABC-based posterior distributions (different colours for different thresholds $\epsilon$) are presented against the true posterior (black). In (a)-(c) we see a reasonable performance of ABC, though clearly less accurate than in the case of linear system. In (d)-(f), however, the algorithm fails to reproduce the posterior in all the cases, resulting in low fidelity (averaged over 10 ABC implementations) within each inset. As expected from Fig. 5(c), the total time $t_{N}$ can now no longer be considered as a sufficient summary statistic, especially around $\Delta=\Delta_{n=2}$.
@@ V.3 Linear optomechanical system
V.3.1 Summary statistic: total time
In Sec. IV.3.2, it was shown that the intensity of emitted light, i.e. the exponent of the exponential waiting-time distribution, should be almost completely characteristic of the emission profile for a linear optomechanical system. Hence, the total time for a fixed number of photon emissions, $t_{N}$ in Eq. (42), should again be sufficient as a summary statistic. However, as the simulation of dynamics is more computationally exhaustive now, we restrict ourselves to 50 photons, i.e. $S(D_{t})=t_{N=50}(D_{t})$. This is (at least approximately) sufficient due to the simplicity of the exponential average waiting-time distribution—assured by the (almost) Poissonian emission profile discussed in Sec. IV.3.2.
Results
@@
We again quantify the performance of the algorithm through the fidelity (46). However, unlike the two-level atom, obtaining true posteriors requires numerical solutions to the dynamics, which are extremely time consuming. We instead show the results of three indicative trajectories for $\Delta=-1\kappa$ in Fig. 7(a)-(c) and for $\Delta=-7\kappa$ in Fig. 7(d)-(f), while the insets now show just the fidelity averaged over 10 implementations of the ABC algorithm—being no longer averaged over different photon-click trajectories. Nevertheless, we see reasonable agreement for all the trajectories considered, suggesting the performance is relatively consistent.
In particular, we also see consistent performance for the different values of detuning. This is in agreement with expectations from the SNR $\gamma$ presented in the inset of Fig. 5(b), which is mostly invariant to changes of the detuning. Moreover, as the statistics form an exponential distribution, the emitted light is almost coherent with no inter-photon correlations, $\forall\tau:\;g^{(2)}(\tau)\approx 1$ in Fig. 4, meaning that the total time for a fixed number of emissions (42) is highly informative. We therefore see an improved performance compared to the two-level atom.
@@ V.4 Non-linear optomechanical system
Finally, we consider the non-linear optomechanical system. As already emphasised, the dynamics is now complex, resulting in non-trivial photon statistics. Nevertheless, we first proceed with the same summary statistics of the total emission time, in order to explicitly demonstrate that, as expected, it will not be satisfactory, especially for detunings yielding photon-bunching effects.
V.4.1 Summary statistic: total time
@@
Results
We show the results compared to true posteriors obtained by numerical means in Fig. 8, as for the linear optomechanical system. For Fig. 8(a)-(c) we set $\Delta=\Delta_{n=1}$, while for Fig. 8(d)-(f) we set $\Delta=\Delta_{n=2}$. As expected, in the $n=1$ regime, we see a good performance of ABC, capable of identifying the parameter range around $\Delta_{n=1}$ to a high degree of accuracy, while mistaking it slightly for $\Delta_{n=3}=-8.4\kappa$ at which the most similar total emission times can occur—recall Fig. 5(c). For the $n=2$ regime, however, the algorithm does not reproduce the posterior, being capable only to eliminate the “inferable” values of the detuning around $\Delta_{n=1}$ from the distribution.
@@ Results
Figure 9: Histogram of waiting times $\tau$ in between subsequent photon-clicks for detuning $\Delta_{n=1}$ in (a) and $\Delta_{n=2}$ in (b). In each case, the data represents average occupancies for 2000 trajectories, each of $N=80$ photo-clicks. By focussing on the short-time regime in the above ($\tau\leq 8\kappa^{-1}$), we see a clear difference between the distributions: (a) – an exponential decreasing modestly due to uncorrelated single-photon ($n=1$) emissions being dominant; versus (b) – a distribution dropping steeply at low $\tau$ due to double-photon ($n=2$) emissions being favoured. The error bars mark the standard deviation of the trajectory-samples used.
@@
Figure 10: ABC performance for the non-linear optomechanical system employing the (truncated) histogram of waiting times as the summary statistic. In particular, the acceptance criterion of the ABC utilises the L2-norm (47) between the histograms representing the waiting-time distribution in the range of short-time intervals, with each trajectory consisting of $N=80$ photon-clicks. As in Fig. 8, two different values of detuning are considered: $\Delta=\Delta_{n=1}\approx-2.8\kappa$ (a-c) and $\Delta=\Delta_{n=2}\approx-5.6\kappa$ (d-f), with each subplot (a-f) being computed for a different photon-click trajectory. Importantly, in all the cases the ABC algorithms is now capable of reproducing the true posterior (black), while a less stringent $\epsilon$-threshold (different colours) must now be chosen due to much higher numbers of samples, $\nu$, being required for the method to converge.
V.4.2 Summary statistic: histogram of waiting times
As confirmed, the total emission time does not constitute a sufficient summary statistic for non-linear optomechanics, as it is monotonic only within restricted ranges of the detuning parameter $\Delta$, see Fig. 5(c), but it is also incapable of distinguishing distinct types of inter-photon correlations. However, before proposing other summary statistics, one could consider a “full-data” approach [66] with the acceptance rule (7) being based directly on the comparison between the observed $D_{t}$ and sampled $D^{\prime}_{t}$ photon-click trajectories. Unfortunately, Eq. (7) would then need to resort to a temporal distance measure [81], $\delta(D_{t},D^{\prime}_{t})$, applicable to point processes [79], which are known, e.g. the Wasserstein distance [82], to be computationally hard, not to mention calibrating then the $\epsilon$-threshold that works within the whole range of $\Delta$ considered. Moreover, the library used within the ABC algorithm would then need to contain the actual trajectories $D^{\prime}_{t}$, as storing summary statistics $S(D^{\prime}_{t})$ labelled by $\Delta$-parameter values would no longer be enough.
Because of these challenges, we propose to use the (truncated) histogram of waiting times in between subsequent photon clicks as a summary statistic, which is known to capture the photon bunching and anti-bunching effects [78, 76] and therefore should be sufficient to discriminate detuning values within the range covering values $\Delta_{n=1}$ and $\Delta_{n=2}$ marked within Fig. 5(c). This is confirmed by Fig. 9, in which we present the so-obtained waiting-time distributions built upon considering 2000 simulated trajectories, each one containing $N=80$ photon-clicks, for the two detuning values of interest: $\Delta_{n=1}$ and $\Delta_{n=2}$.
Although a histogram of waiting times does not contain the full information about a given trajectory—in particular, the ordering of the clicks is lost, so that we expect it to be insufficient around $\Delta_{n=3}$ with photons being then emitted in triplets—we show that it suffices for our purposes. What is more, we ignore the non-stationarity of the underlying optomechanical dynamics, as it occurs at negligible timescales when compared to the total duration times of the trajectories considered.
Importantly, when constructing the histogram we restrict ourselves to inter-click intervals of short duration, i.e. $\tau<8\kappa^{-1}$ in Fig. 9, as we expect the photon bunching effects to alter the—otherwise exponential—waiting-time distribution only at short timescales. Although we lose in this way the information about the total emission time that can be inferred only from the full histogram, we presume it to be (mostly) irrelevant and decide not to include (again) the total time as a secondary summary statistic.
Finally, within the ABC algorithm 1 we employ the L2-norm within the acceptance rule (7), i.e.:
$\delta(S(D_{t}),S(D^{\prime}_{t}))=$ $\sqrt{\sum\limits_{i}\left(x_{i}-x^{\prime}_{i}\right)^{2}}\,,$ (47)
where $S(D_{t})$ is the (truncated) time-binned histogram of the inter-click waiting times as in Fig. 9, and the $x_{i}$’s are the respective sizes of histogram entries.
Results
@@
In Fig. 10, we present the resulting ABC-based posterior distributions, which should be compared directly against the previous ones of Fig. 8 with the same photon-click trajectories being considered. Crucially, the ABC algorithm is now capable of reproducing the true posterior in both the $n=1$ and $n=2$ regimes. For the $n=1$ regime, we see in Fig. 10(a-c) better performance of the ABC than in Fig. 8(a-c) as the values of $\Delta$ around $n=3$ are now excluded within the posterior. However, the distribution itself is no longer that smooth. The latter fact is a consequence of the discretisation induced by the finite width of bins used to construct the histogram and a low number of clicks ($N=80$) composing single trajectories for which the L2-norm (47) is evaluated. Still, despite similar discontinuities we observe for $n=2$ a significant improvement in Fig. 10(a-f). As we tighten the $\epsilon$-thresholds in Fig. 10, we observe that none of the samples is sometimes accepted—indicating that these would require a much higher number of samples than $\nu=4\times 10^{5}$ here used. Nevertheless, for looser thresholds we manage to obtain good results even with the number of samples at hand.
@@ VI Conclusions
In this paper, we have approached the problem of performing quantum Bayesian inference when the underlying quantum system does not allow for efficient computation of the likelihood function, but the simulation of measurement data still remains tractable thanks to the quantum Monte-Carlo techniques [47, 48, 49, 50, 51]. We argue that the reconstruction of the posterior distribution is then still possible by resorting to approximate sampling-based methods. In particular, upon building a sufficiently large library of measurement data, we demonstrate that likelihood-free approaches—in particular, the Approximate Bayesian Computation (ABC) algorithm [61, 62]—can then be used to reconstruct the posterior.
@@
In order to show the applicability of our methods, we consider examples of a two-level atom and an optomechanical system, both being driven by classical light and continuously probed by photodetection. The former allows us to study the trade-off between the accuracy and efficiency of the ABC algorithm, thanks to the inference problem being actually solvable analytically. In stark contrast, in the latter case the sampling-based method is often essential to obtain results within a reasonable computation time. Yet, the understanding of the quantum statistics exhibited by the emitted photons [76] (bunching/antibunching) is then crucial to choose appropriate statistical tests employed within the ABC algorithm allowing it to distinguish well between different datasets.
@@
In each case we demonstrate explicitly how the ABC algorithm should be tailored, in particular, we identify the summary statistic of the measured data that is sufficient to characterise the underlying process. Moreover, we study how to choose the thresholds for the acceptance criteria used in rejection sampling, which, if chosen too tight, may require impractically large numbers of samples. For the non-linear optomechanical system we demonstrate that the histogram of waiting times can be used as a summary statistic that is sensitive to photon-bunching effects, in order for the ABC to reconstruct the posterior distribution reasonably well despite minimal calibration. However, let us note that if one (albeit unlikely) requires to reconstruct the posterior in the full parameter range, then ABC can be used to only infer an “intermediate” posterior that is already narrowed down to relevant parameter ranges based on the data. As a result, such a narrowed posterior can be used afterwards as a prior by more computationally demanding methods, which can then be already effective thanks to the restricted parameter range.
Apart from the open quantum systems considered here, we emphasise again that the methods presented may be applicable to a wide range of physical systems. More precisely, any system that possesses the asymmetry such that it is easy to simulate and sample from its measurement data, but hard to compute the likelihood for, will benefit from the algorithmic approach described in this work. This is the case, for instance, in Hamiltonian learning [64, 65], but one may also think of other scenarios in which combining quantum Monte-Carlo techniques [47, 48, 49, 50, 51] with other methods deems sampling easy while the likelihood computation remains hard. For example, it is so for many-body systems that can be mapped onto spin models [83], so that the quantum-jump dynamics can then be efficiently simulated with help of matrix-product-state (MPS) representations [84, 85, 86].
@@
For the above reasons, we believe that our results pave the way for using likelihood-free inference methods across different experimental platforms. A natural next step is to verify their performance in particular quantum estimation tasks, while combining them with other statistical inference techniques such as the Markov chain Monte Carlo methods and the Metropolis-Hastings algorithm, as well as particle filters [54]. These are widely used to avoid the need of considering the full parameter space that, in contrast to our work, may pose a serious challenge when inferring multiple (multidimensional) parameters from the measurement data. We leave this as an interesting line of research to be pursued in the future.